### v1

# StackGen: Generating Stable Structures from Silhouettes via Diffusion

Abstract
Humans naturally obtain intuition about the interactions between and the stability of rigid objects by observing and interacting with the world. It is this intuition that governs the way in which we regularly configure objects in our environment, allowing us to build complex structures from simple, everyday objects. Robotic agents, on the other hand, traditionally require an explicit model of the world that includes the detailed geometry of each object and an analytical model of the environment dynamics, which are difficult to scale and preclude generalization. Instead, robots would benefit from an awareness of intuitive physics that enables them to similarly reason over the stable interaction of objects in their environment. Towards that goal, we propose StackGen—a diffusion model that generates diverse stable configurations of building blocks matching a target silhouette. To demonstrate the capability of the method, we evaluate it in a simulated environment and deploy it in the real setting using a robotic arm to assemble structures generated by the model. Our code is available at https://ripl.github.io/StackGen.
I Introduction
Understanding the physics of a scene is a prerequisite for performing many physical tasks, such as stacking, (dis)assembling, and moving objects. Humans can intuitively assess and predict the stability of structures through a combination of visual cues, force feedback, and experiential knowledge. On the other hand, robots lack natural multimodal sensory integration and an understanding of intuitive physics. Robots have traditionally relied upon a world model that includes a representation of the detailed geometry of the objects in the environment and an analytical model of the dynamics that govern their interactions. This dependency poses significant challenges to deploying robotic agents in unprepared environments.
The ability to compose a diverse array of blocks into a stable structure has a long history as a testbed to study an agent’s understanding of object composition and interaction [1, 2, 3, 4]. While seemingly primitive, this ability comes with many practical implications such as robot-assisted construction [5, 6, 7, 8, 9], and would serve as a backbone for downstream applications where an agent deals with complex sets of real world objects.
Contemporary approaches to building 3D structures based upon an intuitive understanding of physics utilize the predicted forward dynamics of a scene as part of a planner that combines building blocks into a target structure. This typically involves first training a forward dynamics model that serves as the intuitive physics engine, and then using this model to simulate the behavior of candidate object placements via a form of rejection sampling. Such an approach comes at a high cost as it requires searching through a large space of coordinates and modeling the dynamics for each possible block placement.
Figure 1: StackGen consists of a diffusion model that takes as inputs a target structure silhouette and a list of available block shapes. The model then generates a set of block poses $\hat{\bm{p}}_{1},\ldots\hat{\bm{p}}_{k}$ that construct a stable structure consistent with the target silhouette. The resulting structure can then be constructed using a robot arm.
Rather than training a forward dynamics model, we consider learning and generating a joint distribution over the SE(3) poses of objects composed to achieve a stable 3D structure (Fig. 1). We condition this distribution on a user-provided specification of the structure, allowing them to control the generation at test time.
Inspired by its success in computer vision [10, 11, 12, 13, 14] and, more recently, robotics domains [15, 16, 17], we employ conditional diffusion models [18], a family of generative models shown to perform well in various generation domains, in our case for producing stable 6-DoF object poses. Similar in spirit with those approaches that control image generation via spatial information such as sketch or contour [19], we ask a user to provide a silhouette that vaguely describes the desired structure, and use it as a conditioning signal. Different from Zhang et al. [19], we simply train a conditional diffusion model built on the Transformer architecture. We should note that, unlike standard image generation, our approach generates a set of poses. And we aim to generate those that result in a physically stable structure.
Our model (StackGen) reasons over the 6-DoF pose of different building blocks to realize their composition as part of a stable 3D structure consistent with different user-provided target specifications. In the following sections, we describe a transformer-based architecture that underlies our diffusion model and the procedure by which we generate stable block configurations for training and evaluation. We evaluate the capabilities of StackGen through baseline comparisons and as well as real-world experiments that demonstrate its benefits to real-world scene generation with a UR5 arm.
II Related Work
II-A Learning Stability from Intuitive Physics
Similar to our work, several efforts [2, 3, 15] consider the interaction of relatively simple objects to investigate the notion of intuitive physics. When considering visual signals for assessing stability, ShapeStacks [20] successfully learned the physics of convex objects in a single-stranded stacking scenario. This is achieved by vertically stacking objects, calculating their center of mass (CoM), and scaling up the dataset to train a visual model via supervised learning, enabling stability prediction prior to stacking. However, calculating the CoM for combinations in multi-stranded stacks proves to be much less straightforward. Another form of intuitive physics involves the ability to predict how the state of a set of objects will evolve in time, which includes concept of continuity and object permanence. This has motivated the development of benchmarks that measuring models’ ability on such tasks called violation-of-expectation (VoE) [21, 22, 23, 24]. In the context of robotics, Agrawal et al. [25] collect video sequences of objects being poked with a robot arm. Using this dataset, they train forward and inverse dynamics models from pixel input and demonstrate that the model enables the robot to reason over an appropriate sequence of pokes to achieve a goal image. Other work has similarly followed suit [26, 27].
II-B Diffusion Models for Pose Generation
Given their impressive ability to learn multimodal distributions, a number of works employ diffusion models [28] to learn distributions over the SE(3) poses in support of robot planning [29, 30, 31]. Urain et al. [29] use conditional diffusion models to predict plausible end-effector positions conditioned on target object shapes for robot manipulation. Simeonov et al. [30] use a diffusion model to predict the optimal placements of objects in a scene by modeling the spatial relationships between objects and their environment, identifying target poses for tasks like shelving, stacking, or hanging. Their method incorporates 3D point cloud reconstruction as contextual information to ensure that the predicted poses are both functional and feasible in real-world scenarios. Liu et al. [32] and Xu et al. [33] combine large language models with a compositional diffusion model to analyze user instructions and generate a graph-based representation of desired object placements. They then predict object arrangement patterns by optimizing a joint objective, effectively merging language understanding with spatial reasoning.
II-C Automated Sequential Assembly
Relevant to our data generation procedure, Tian et al. [34] propose an assembly method (ASAP) that relies on a reverse process of disassembly, where each component is placed in a unique position to guarantee physical feasibility. However, this approach does not account for the potential structural instability that might arise from multiple combinations, since the assembly scenario assumes a one-to-one mapping of components to specific locations.
In contrast, our work addresses the challenge of finding a structure that maintains gravitational stability using only a 2D silhouette through a one-to-many mapping approach. This method ensures that the structural stability is retained and accurately reproduced when transitioning to a 3D environment. Our focus is on the generation and verification of structurally stable block configurations rather than optimizing the assembly sequence. Similar to the method of ASAP, which generates step-by-step assembly sequences where intermediate configurations remain stable under gravitational forces, we propose a “construction by deconstruction” method that enables scalable data generation by predicting diverse stable configurations, without relying on predefined assembly paths.
III Method
Figure 2: A visualization of StackGen’s transformer-based architecture.
In this section, we describe our diffusion-based framework for generating SE(3) poses for blocks that together form a stable structure consistent with a user-provided specification of the scene. We then discuss the procedure for training the model, including an approach to producing a training set that contains a diverse set of stable block configurations.
III-A Diffusion Models for SE(3) Block Pose Generation
Our model (Fig. 2) generates the SE(3) block poses necessary to create a 3D structure that both matches a given condition (e.g., a silhouette) and is stable. Underlying our framework is a transformer-based diffusion model that represents the distribution over stable 6DoF poses, without explicitly specifying the number, type, or position of its constituent blocks. In this way, the model employs a reverse diffusion process to produce block poses that collectively form a stable structure. Separately, we train a convolutional neural network (CNN) to predict the number and type of blocks necessary for the construction based on the target silhouette. At test-time, we employ the CNN to predict the block list, and then provide this list and the target silhouette as input to the diffusion model. The diffusion model then samples potential block poses composing a stable structure.
We adopt denoised diffusion probabilistic models (DDPM) [28] as the core framework of our model. Following DDPM, we start with a forward diffusion process that adds noise to the state space of interest. In our case, given a set of block poses $\bm{p}_{1},\bm{p}_{2},\ldots,\bm{p}_{k}\in\mathbb{R}^{d}$ that compose a stable stack and a diffusion timestep $t\in[1,T]$ that specifies a noise scale. [^1] We normalize poses before applying the diffusion framework. During inference, the generated poses are unnormalized accordingly. Diffusion training with noise-injection follows as:
$\tilde{\bm{p}}^{t}_{i}=\sqrt{\bar{\alpha}_{t}}\bm{p}_{i}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{i},~{}~{}\bm{\epsilon}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $\bar{\alpha}_{t}$ is a coefficient determined by a noise schedule and $\bm{\epsilon}_{i}$ is fresh noise injected in each step. StackGen then provides the noisy poses $\tilde{\bm{p}}^{t}_{i}~{}(i=1,2,\ldots,k)$ to our denoising network $D_{\theta}$ along with the diffusion timestep $t$, shapes $s_{1},\ldots,s_{k}$ and the silhouette of the blocks $S$. This results in the expression
$\hat{\bm{\epsilon}}_{1:k}=D_{\theta}(\tilde{\bm{p}}_{1:k},t,s_{1:k},S),$ (2)
where the notation $X_{1:k}$ is equivalent to $X_{1},\ldots,X_{k}$, and $\bm{\epsilon}_{i}$ is a predicted pose noise for the $i$-th block. With the predicted noises, the training objective for a single sample is
$\frac{1}{k}\sum_{i=1}^{k}\|\bm{\epsilon}_{i}-\hat{\bm{\epsilon}}_{i}\|^{2}.$ (3)
We sample diffusion timestep $t$ uniformly random from $[1,T]$ at each training step.
Once the denoising network is trained, the sampling procedure starts with sampling a noisy pose from Gaussian distribution
$\tilde{\bm{p}}^{T}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$
From this initial noises, we iterate the following step from $t=T$ to $1$
$\tilde{\bm{p}}^{t-1}_{i}=\frac{1}{\sqrt{\alpha}_{t}}\left(\tilde{\bm{p}}^{t}_{i}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\bm{\epsilon}}_{i}\right)+\sigma_{t}\bm{z},$ (4)
where $\bm{\epsilon}_{i}$ is given by Eq. 2 and $\bm{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ if $t>1$, otherwise $\bm{z}=\bm{0}$. The resulting $\tilde{\bm{p}}^{0}_{1:k}$ are the generated poses.
III-B Model Architecture
Challenging requirements for the model come from the nature of the task that 1) the model must be able to work with a variable number of block poses since different stacks use different number and shapes of blocks; and that 2) the model must process inputs from different modalities, including poses, shapes and silhouette that has spatial information.
Our model (Fig. 2) is built upon the Transformer architecture [35], which can process input tokens that may originate from different modalities. To initialize the process, we use a convolutional neural network (CNN) to predict the block list of a structure from that structure’s silhouette, shown in Figure 1. For training we uniquely encode the number of cubes, rectangles, long rectangles and triangles in a structure using an integer index and proceed by training the CNN $C_{\theta}$ with parametrization $\theta$ to model the joint distribution of block counts, represented by the class probability of each index, using the cross-entropy loss
$L(D,\theta)=\frac{1}{|D|}\sum_{(S_{i},y_{i})\in D}-\log\frac{\exp(C_{\theta}(y_{i}|S_{i}))}{\sum_{y_{k}}\exp(C_{\theta}(y_{k}|S_{i}))},$ (5)
where $D=\{(S_{i},y_{i})\}$ is our labeled training set, $S_{i}$ is the structure’s silhouette, and $y_{i}$ is the index corresponding to the structure’s block list. This predicted block list serves as one of the inputs to the subsequent steps of our model.
Given a scene that contains stable stack of $k~{}(\leq N)$ blocks, we extract a list of their poses $\bm{p}_{1},\bm{p}_{2},\ldots,\bm{p}_{k}\in\mathbb{R}^{6}$ and shape embeddings $\bm{s}_{1},\bm{s}_{2},\ldots,\bm{s}_{k}\in\mathbb{R}^{d}$. We use a 6-dimensional pose representation consisting of Cartesian coordinates for translation and exponential coordinates for orientation. The shape embedding is retrieved from a codebook storing unique trainable embeddings for each shape, and the poses are projected into $\mathbb{R}^{d}$ space with an MLP applied independently to each pose. A diffusion timestep $t\in[1,T]$ is also converted to an embedding $\bm{t}\in\mathbb{R}^{d}$. The pose, shape, and diffusion timestep embeddings are summed for each object to obtain $k$ object tokens.
To handle variability in the number of blocks, we make the number of input tokens to the Transformer encoder constant by padding the remaining $N-k$ object tokens with zero vectors, resulting in $N$ tokens independent of $k$.
The silhouette of the block structure is given as a binary image of size $64\times 64$. Following Dosovitskiy et al. [36], we split this into $16$ patches of $16\times 16$ each and use a two-layer MLP to encode each patch independently to obtain silhouette tokens. Sinusoidal positional embeddings [35] are added to the silhouette tokens to retain spatial information.
The object and silhouette tokens are then combined and fed into the Transformer encoder. At the last layer of the encoder, each contextualized block token is projected back to pose space ($\mathbb{R}^{6}$) and supervised with the original noise added to the corresponding pose, following the framework of DDPM. Figure 2 summarizes this process and architecture.
III-C Generating Data
Figure 3: Our strategy (left) to generate diverse set of stable stacks. After filling the design grid with shapes, we verify stack stability in simulator, and begin removing each block, saving the stable stacks. The right part shows some challenging examples in the dataset.
To train a model that can generate diverse set of stable block poses, the quality and diversity of the dataset is crucial. We seek to have an algorithm that synthetically samples various stable block configurations to generate such dataset at scale. If we place excessive emphasis on diversity of the block stacks, a naive and general approach could be to spawn and drop a randomly selected shape at a random pose in simulation, wait until it settles and repeat this process until a meaningful stack gets constructed in the scene (by checking their height or collisions between blocks, for example). In the case that this process does not end up in a stack, we could reject it and start over again, repeating the procedure. This could potentially lead to a very general and extremely diverse dataset of block stacking, however, it was found to be inefficient and impractical.
As an alternative, we employ a “construction by deconstruction” approach that involves starting with a dense structure comprised of different block shapes, followed by a block removal process that involves iteratively removing blocks from the stack until it becomes unstable. While the initial structure is guided by a pre-defined grid, we find that the random horizontal displacement and block removal process creates a diverse set of non-trivial structures.
[…]
For each stack of blocks generated as above, we proceed to generate additional data points via block removal, whereby we remove blocks whose absence does not collapse the structure. From the initial stack of blocks, as depicted in Figure 3, we try removing each block and simulate the effect on the remaining blocks in the stack. If the stack remains stable, we add the resulting set of block poses and the silhouette to the dataset, and then repeat with another block. We apply this procedure recursively to each stable configuration, removing at most four blocks. We note that the block at the top of the stack is excluded from removal, and thus data samples always have the height of four cubes. Following this procedure, we generate $191$k instances of stable block stacks that we then split into training and test sets using a 9:1 ratio.
IV Experiments
We evaluate the ability of our model to generate a stable configuration of objects that is consistent with a reference input that can take the form of an example of the block structure or a sketch of the desired structure (Fig. 1). We then present real-world results that involve building different structures using a UR5 robot arm.
IV-A Evaluation in simulation
We evaluate our model using a held-out test dataset. Figure 5 shows generated stacks paired with their corresponding silhouettes. In Figure 7, we present a diverse set of stacks produced by the model for a single silhouette, demonstrating its capability for multimodal distribution learning.
Figure 5: Silhouettes from the heldout dataset and rendering of block poses generated by our model.
We aim to evaluate our approach with two metrics: 1) the frequency with which our method generates block poses that form a stable structure; and 2) the consistency of the generated stacks with the target silhouette. The problem of generating a stable structure from a given block list and silhouette can have multiple solutions, so our evaluation technique samples three sets of block poses for each pair of silhouette and block list in the test set. We compare our method against two baselines, the Brute-Force Baseline and the Greedy-Random Baseline.
IV-A1 Brute-Force Baseline
Given a silhouette and a set of available blocks, this algorithm searches for potential placement poses of each block by maximizing a silhouette alignment score given by silhouette intersection while minimizing a collision penalty between predicted blocks. To achieve a high alignment score, for each block we sample 20 coordinates ${(x_{i},0,z_{i})}$ where $x_{i}$ and $z_{i}$ are sampled uniformly from $[-3,3]$ and $\{1,3,5,7\}$, respectively. We perform 20 linear searches from each point along the x-axis in both directions to find poses with optimal alignment and collision measures.
IV-A2 Greedy-Random Baseline
This approach uses a left-to-right, bottom-to-top algorithm operating on the structure silhouette to place blocks. Starting from the lowest layer to the highest, each layer is assigned a fixed height. The algorithm measures the distance of the longest consecutive line of pixels from left to right. It then considers all blocks in the current block list whose width is less than this distance and greedily places the longest one. Since this algorithm is deterministic, we introduce a swap mechanism to add diversity: with a certain probability $\sigma$, the algorithm will swap two adjacent cubes within the same layer with a rectangle elsewhere in the silhouette (since two cubes and a rectangle are of equal length). By controlling the probability $\sigma$, we can adjust the diversity of the generated configurations. This swap mechanism is also applied to the Brute-Force baseline to control its variability.
To quantify the diversity of predicted poses, we analyze the predicted results by obtaining a per-layer block list arranged from left to right. Two constructions are considered distinct if their layered block lists differ. The diversity metric is then defined as the number of distinct poses sample poses generated for a given input divided by the total number of samples taken. Figure 7 contains two scenes with average diversity of 83.33%.
TABLE I: Three Side View IoU (%)
Front Side Above Average
Brute-Force Baseline 63.11 58.04 54.49 58.55
Greedy-Random Baseline 58.45 54.50 54.19 55.72
StackGen (ours) 77.03 76.36 70.47 74.62
For stability evaluation, we spawn blocks according to their generated poses, observe their subsequent behavior (i.e., using a simulator for non-real-world experiments), and check whether any of the blocks fall to a layer below where they began, which would lead the sample to be classified as unstable. To assess silhouette consistency, we extract the silhouette of the generated structure after running forward dynamics, compute the intersection over union (IoU) for the silhouettes from three different views (front, side, and top, shown in Table I), and then calculate the average IoU across these views. Unstable (collapsed) structures receive an IoU of zero.
Figure 7: A (left) reference (ground-truth) stack with its silhouette, and (right) a diverse set of structures generated from the silhouette by our model.
TABLE II: Stability and Consistency with Same Diversity
Stability (%) $\uparrow$ IoU (%) $\uparrow$
Brute-Force Baseline 68.13 (1022 / 1500) 58.55
Greedy-Random Baseline 71.93 (1079 / 1500) 55.72
StackGen (Ours) 86.67 (1300 / 1500) 74.62
We evaluated all three models using our pretrained CNN block list predictor on 500 scenes, with three samples generated per scene. Since StackGen achieved a diversity level of $60.47\%$, we set $\sigma=0.6$ in the Greedy-Random Baseline to match this diversity level. As shown in Table II, StackGen significantly outperforms the baselines in both stability and IoU.
IV-B CNN Ablation: Predicted vs. Ground-Truth Block Lists
To ensure that our CNN model did not skew the overall results, we conducted an ablation study following the main experiments. Using 500 scenes from the test dataset, we generated three samples per scene, applying both the CNN-predicted block list $\{\hat{S}_{1:k}\}$ and the ground-truth block list $\{S_{1:k}\}$. As shown in Table III, the quantitative results showed no more than a 2% difference in stability or IoU. These findings confirm that the CNN does not introduce any bottlenecks in our framework. Therefore, all other comparisons are conducted using our CNN predictor.
TABLE III: Ground-Truth vs. Predicted Block List
Stability (%) IoU (%)
StackGen (w/ GT) 88.13 (1322 / 1500) 76.21
StackGen (w/ CNN) 86.67 (1300 / 1500) 74.62
IV-C Block Stacking in the Real World
[…]
Figure 8: Examples of various stable 3D structures constructed by a UR5 robot arm based upon goal specifications in the form of 8(a) images and 8(b) sketches of the target structure. Note that StackGen seeks to match the silhouette of the input and as a result, the color and type of individual blocks may differ from the reference input.
To demonstrate that our method performs well in a real-world environment, we conducted an experiment using toy blocks and a UR5 robotic arm. Our goal was to build a pipeline that operates as follows: first, a user provides a silhouette by either presenting a reference stack of toy blocks or drawing a sketch of their desired structure. After extracting a silhouette from the stack or sketch, our model generates a stable configuration of blocks that matches the provided silhouette. Finally, the UR5 arm assembles the generated stack on a table using real blocks.
[…]
In this scenario, a silhouette is extracted from a stack using a simple rig consisting of an RGBD camera (Realsense 435D), toy blocks, and a white background, as shown in Figure 1. The rig captures a photo of a stack of blocks built by the user then makes a binary silhouette by filtering out background pixels using depth readings and applying a median filter to smooth the silhouette, removing any remaining white pixels, finally resizing and pasting the result onto a $64\times 64$ canvas.
[…]
In this case, we use the camera to capture a hand-drawn sketch from a user (Figure 1). It is converted into a binary image, smoothed using a median filter, and a bounding box with a $4\times 4$ grid is put around it. We then compute the occupancy to identify whether each grid cell is fully or partially occupied (e.g., a triangle).
[…]
Out of the eight cases we tested, this pipeline successfully built all of the stacks stably, with only minor discrepancies relative to the original silhouettes (considering the error of block initial position). However, we note that this does not imply that our system is flawless. As discussed in Section IV-A, the model can sometimes generate unstable block configurations. Nonetheless, in these real-world experiments, the success rate indicates that the model is robust enough to handle potentially out-of-distribution silhouettes effectively.
V Conclusion
In this paper, we presented a new approach that enables robots to reason over the 6-DoF pose of objects to realize a stable 3D structure. Given a dataset of stable structures, StackGen learns a distribution over the SE(3) pose of different object primitives, conditioned on a user-provided silhouette of the desired structure. At inference time, StackGen generates a diverse set of candidate compositions that align with the silhouette while ensuring physical feasibility.
We conducted experiments in a simulated environment and showed that our approach effectively generates stable structures following a user-provided silhouette, without modeling physics explicitly. Further, we deployed our approach in a real-world setting, demonstrating that the method effectively and reliably generates stable and valid block structures in a data-driven manner, bridging the gap between visual design inputs and physical construction.